Linda wants to open a T-shirt stand for Homecoming week. The school will license her a booth for $100. Each T-shirt will cost her $4. Linda's average total cost function will be:

Select one:

a. $104 / X

b. $100 + $4 * X

c. $100 / X + $4 / X

d. $100 / X + $4
Answer #1

Ans) the correct option is d. $100 / X + $4

Total cost = 100 + 4*X

Average total cost = total cost / quantity =  (100 + 4*X)/ X = 100/X+ 4